OpenStudy (philcua):

when you multiply 2 negative fractions is the answer gonna be positive

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

yes negative time negative = positive

OpenStudy (anonymous):

\[-\frac{a}{b} \times - \frac{c}{d} = \frac{ac}{bd}\]

OpenStudy (s117):



OpenStudy (philcua):

thx guys thx so much wish i could give all medal

OpenStudy (anonymous):

no prob

OpenStudy (philcua):

anyone here play calsh of clans